OpenStudy (anonymous):

Find the y-intercept. 5 + y = -3x y = -3x - 5 y-intercept = __

OpenStudy (smored):

y=mx+b y-intercept = b

OpenStudy (anonymous):

i don't get it..wut

OpenStudy (ipwnbunnies):

He gave you the general form of the slope-intercept equation. The 'b' in the general form is the y intercept. Breh.

OpenStudy (anonymous):

the y intercept when x is 0 therefore y=b -3(0)=0